OpenStudy (anonymous):

Can someone please explain this like really break it down for me that even my two year old niece can understand. ... 4x-26-3x=16-20 I'm supposed to find the solution to the equation.

OpenStudy (phi):

on the right side you have 16-20 any idea how to simplify that ?

OpenStudy (anonymous):

You mean make it smaller?

OpenStudy (phi):

yes, do the subtraction

OpenStudy (anonymous):

Ok 16- 20 = - 4

OpenStudy (phi):

so 4x-26-3x=16-20 becomes 4x-26-3x= -4 that is progress. on the left side you have 4 x's take away 3 x's how many x's do you have ?

OpenStudy (anonymous):

7

OpenStudy (phi):

if you have 4 spoons and take away 3 spoons you don't have 7 spoons if you have 4 x's and take away 3 x's you don't end up with 7 x's

OpenStudy (anonymous):

sorry I didn't mean to add and I was using my phone as old as I am I should know this the answer is 3

OpenStudy (phi):

4x-3x = 1x we don't usually bother to write 1x (which means 1 times x), we just write x so 4x-3x= x

OpenStudy (phi):

that means 4x-26-3x= -4 can be rewritten as x - 26 = -4 now add +26 to both sides, like this: x+26-26 = +26-4 simplify both sides (do the subtraction)

OpenStudy (anonymous):

yeah I was okay once I found out 4x-3x=1 or just x but after that I got lost if I cant learn this how can I help my own son with his math homework? do you want me to add 26 to 1 then minus 26 ??

OpenStudy (phi):

we are at x -26= -4 we can add the same thing to both sides, and things will stay equal x-26+26= -4+26 you can add -26 and +26 on the left side but x is not a "simple" number. in other words, x is not the same as 1. It is x but we can simplify the pure numbers can you try to do that and write down the new equation ?

OpenStudy (anonymous):

going to watch them now then ill come back thank you